# Nagaoka Ferromagnetism in $3\times 3$ Arrays and Beyond

### v1

@@ Abstract
Nagaoka ferromagnetism (NF) is a long-predicted example of itinerant ferromagnetism (IF) in the Hubbard model that has been studied theoretically for many years. The condition for NF, an infinite onsite Coulomb repulsion and a single hole in a half-filled band, does not arise naturally in materials, and was only realized recently for the first time in experiments on a $2\times 2$ array of gated quantum dots. Atomically precise fabrication of dopant arrays in Si allows for engineering highly controllable systems with complex geometries. This makes dopant arrays a good candidate to study NF in different array geometries through analog quantum simulation. Here we present theoretical simulations done for $3\times 3$ arrays and larger $N\times N$ arrays of quantum dots, and predict the emergence of different forms of ferromagnetism in different geometries. We find NF in perfect $3\times 3$ arrays, as well as in $N\times N$ arrays for one hole less than half-filling. The ratio of the Hubbard on-site repulsion $U$ to hopping $t$ that defines the onset of NF increases as N increases, approaching the bulk limit of infinite $U$ for large N. Additional simulations are done for geometries made by removing sites from $N\times N$ arrays. Different forms of ferromagnetism are found for different geometries. Loops show ferromagnetism, but only for three electrons. For loops, the critical $U/t$ for the onset of ferromagnetism decreases as the loop length increases. We show that this different dependence on size for loops and $N\times N$ arrays can be understood by a scaling argument. Our results show how analog quantum simulation with small arrays can elucidate the role of effects including wavefunction connectivity; system geometry, size and symmetry; bulk and edge sites; and kinetic energy in determining the quantum magnetism of small systems.
@@ I Introduction
NF is one of the exact results for itinerant ferromagnetism (IF) and has been studied theoretically for years in various conditions [2, 3]. However, the specific conditions proved by Nagaoka do not arise in any natural materials and has only been observed experimentally for the first time in an engineered $2\times 2$ quantum dot plaquette [4]. It remains an open question when NF can be extended to larger, finite-systems that have internal as well as edge sites, how NF would behave in realistic conditions where $U$ is large but not infinite, and how NF is affected by long-range interactions or disorder.
Following Nagaoka’s original work, many have tried to generalize Nagaoka’s result. Tasaki provided a proof of a generalized version of Nagaoka’s theorem [2]. Mielke and others found a new class of ferromagnetism states in the Hubbard Model which they called “flat-band ferromagnetism” [5]. The possibility of extending NF to finite $U$ with finite densities of holes has also been studied for different infinite lattices [6, 7]. There have also been many theoretical works done on the instability of NF, that show when NF does not take place [8, 9, 10].
@@
Quantum dots, also known as artificial atoms, can be used as quantum simulators which can perform simulations beyond what classical computers can do[15, 16, 17, 18]. Among various experimental platforms, dopant-based quantum dots have many advantages, such as the precise placement of dopant atoms and tunable site-to-site hopping amplitude, making them a strong candidate for simulating strongly correlated systems, like the Hubbard Model in the large $U$ limit. Silver recently fabricated $3\times 3$ arrays of single/few-dopant quantum dots and demonstrated analog quantum simulation of a 2D Fermi-Hubbard Model with tunable parameters [19]. This provides an experimental pathway to probe the NF in $3\times 3$ arrays and in many other finite geometries.
@@
In this paper, we use ED to simulate $3\times 3$ dopant arrays in the conditions required for NF, i.e. one less electron than a half filled band and in the large $U$ limit. We predict that NF will exist in a $3\times 3$ array, as in a $2\times 2$ array, as well as in $N\times N$ arrays, for a finite region of $t/U$, where $t$ is the nearest-neighbor hopping. For $t/U$ below the transition point, the ground state of the system has maximum total spin $S$, i.e. the ground state is ferromagnetic. Since the direction of the ferromagnetism is arbitrary, the system’s energy only depends on total spin $S$ but not the spin in the $z$-direction ($s_{z}$). The ferromagnetic state with maximum $s_{z}=S$ with all of the spins aligned in the same $z$ direction is degenerate with all other ferromagnetic states with smaller $s_{z}$ as long as they have the same total $S$. The ground state is $(2S+1)$-fold degenerate when there is Nagaoka ferromagnetism. Above the transition point, the ground state is no longer the maximum total spin state and is no longer ferromagnetic. We simulate variations of the perfect $3\times 3$ array to determine when NF is robust to experimental impurities or disorder.
@@
The quantum nature of magnetism plays an essential role in determining how and when magnetism occurs. The small-scale simulators that we consider in this paper provide an excellent testbed for investigating quantum effects. We find that the connectivity of the many-body wavefunction plays a key role in determining when NF can occur or when other forms of itinerant ferromagnetism at fillings other than one hole in a half-filled band, that is, rather than NF, can occur. Band filling is another key condition that determines the type of ferromagnetism. We will show that very different fillings, related to the number of holes is a half-filled band or the number of electrons in an empty band will define the possibility of ferromagnetism in similar, related structures. The competition between kinetic energy and coulomb repulsion determines where in $t/U$ space the transition to ferromagnetism occurs. The $t/U$ ratio at the transition can increase or decrease with system size, depending on the array geometry (and the related wavefunction connectivity), and the band filling. We develop a simple scaling argument to explain both types of size-dependence and the critical role that the kinetic energy plays, through its dependence on band filling, in determining the form of the scaling. The role of these quantum effects will be highlighted as we discuss NF in different small arrays to show how quantum simulation with small arrays could reveal these critical effects. It should be emphasize that dopant arrays, once they can be made perfectly [20] would be an versatile solid-state platform to explore the diversity of arrays and the range of effects that we discuss.
@@
The paper is structured as follows: In Section II, we discuss the 2D Hubbard Model, the signatures of finding NF in finite systems, and how the systems are modelled in our simulations. In Section III, we present our simulation results for the $3\times 3$ and $N\times N$ square arrays, which exhibit NF up to a finite $t/U$ transition point that decreases with increasing N. In Section IV, we describe the robustness of NF in a $3\times 3$ array with disorder. Finally in Section V, we summarize and discuss essential results about quantum magnetism that can be revealed by simulations done on small arrays of dopants or quantum dots.
@@ II Hubbard Model
where $t$ is the hopping between nearest neighbor sites, $c^{{\dagger}}$ and $c$ are the creation and annihilation operators, $n$ is the particle number operator, and $U$ is the onsite interaction between spin-up and spin-down electrons. NF is determined by a competition between hopping energy and repulsion, hence the parameter that matters is $t/U$. Some use the hopping term $t$ with a minus sign, but here we use the convention without a minus sign and allow the value of $t$ to be positive or negative. The sign of $t$ is important because in some systems the Nagaoka theorem can be proved for one sign but not the other. The sign of t can often be changed with a basis change. In particular, the transformation between the signs of $t$ can be done for a bipartite system. In these cases, NF, if it occurs, will occur for both signs of $t$. The $N\times N$ arrays we study are bipartite lattices.
To find out if a system has NF, it is necessary to determine whether the one-hole ground state has maximum spin. For a finite system like the $2\times 2$ array with 3 electrons, the Hamiltonian is small and the energy spectrum can be calculated directly, as shown in [4]. The energies for the two different total spins, $S=3/2$ and $S=1/2$ are plotted against the parameter $t/U$ in Fig. 1a. Both energies are shifted with respect to the energy of the maximum-total-spin state $E_{S=3/2}=-2t$. Below $|t/U|\approx 0.053$, the ground state of the system is the maximum-total-spin $S_{max}=3/2$ state. In the ferromagnetic ground state, the states with $S=S_{max}$ have a degeneracy of $12S_{max}+$ 1. In the $2\times 2$ array with 3 electrons, there are four degenerate $s_{z}$ states with $s_{z}=-3/2,-1/2,1/2,3/2$. Because the Hubbard Model does not include spin-mixing, it is convenient for the ED to work in $s_{z}$ subspaces. In our simulation, instead of plotting the ground state energy for each of the total spins, we plot the ground state energy for the different spin-$z$ subspaces as shown in Fig. 1b, and shift them by the energy of the state with maximum $s_{z}$ $E_{s_{z}=3/2}=-2t$. Below the transition when the system is in the NF regime, the ground state is $4$-fold degenerate. The signature of NF is when the ground states of the spin-$z$ subspaces become degenerate with each other. This guarantees that the ground state of the system has maximum total-spin.
@@
Figure 1: Nagaoka ferromagnetism in a $2\times 2$ array. (a) Energy spectrum ($S=3/2$ and $S=1/2$ as a function of $|t/U|$), which are offset by the energy of $S=3/2$. Black dotted line indicates the transition for NF. (b) Ground state for each of the different spin z subspaces as a function of $|t/U|$, which are offset by the energy for $s_{z}=3/2$. Above and below the transition point, the ground state for $s_{z}=\pm 1/2$ switches from the total $S=3/2$ to $S=1/2$.
Fig. 1 shows how energy behaves as $|t/U|$ varies. The lattice is bipartite so the absolute sign of the $t$ does not matter here, and $t$ and $-t$ give the same energies. This is not always the case as seen later when the system is no longer bipartite, and the ±$\pm t$ cases can behave very differently. In Fig. 1b, the $s_{z}=\pm 1/2$ state is always a ground state, but the ground state after the transition has a different total $S$ and degeneracy compared to before the transition. This change in total spin can be verified by direct calculation of the total spin for each spin $z$.
The NF in $2\times 2$ array has been verified experimentally in [4] for a $2\times 2$ gated quantum dot plaquette. An ab initio calculation of NF in quantum dots was done in [21]. Its robustness was discussed in [21] and theoretically in [13].
@@ III.1 $3\times 3$ case
The $2\times 2$ array has no internal site. However, the $3\times 3$ dopant array studied by [19] has one internal site. In this section, we look at $3\times 3$ arrays and beyond to see how NF behaves as the size of the system increases and the system approaches the 2D bulk limit. NF has not yet been studied in these larger systems, which should be accessible with dopant arrays. Our results will define the $t/U$ ratios needed for NF. For the $3\times 3$ array, the ground state energy for each spin-$z$ subspace is calculated to determine the presence of NF, and if it exists, to identify the point at which the transition occurs. The half-filled band with one hole has 8 electrons, and the maximum total spin is 4. If all $9$ $s_{z}$ subspaces have the same ground-state energy, the system has NF.
@@
Figure 2: NF in a $3\times 3$ array. Black dotted line indicates the transition for NF. Ground state for each of the different spin $z$ subspaces plotted as a function of $|t/U|$, which are offset by the energy for $s_{z}=4$. When the plotted energy difference for an $s_{z}$ ground state becomes negative, the total spin of the ground state for that $s_{z}$ switches from $S_{max}=4$ to a lower $S$.
@@
We plot the ground state (GS) energy for each subspace in Fig. 2 (again shifted by the GS energy for maximum $s_{z}$). For a small $t/U$ or a large $U$, the $s_{z}$ states are all degenerate, i.e. the system has a maximum spin ground state. This is in contrast to the usual anti-ferromagnetic spin order near half-filling, where states have increasing energy as $s_{z}$ increases. The maximum spin GS holds up to $|t/U|\approx 0.0144$ (black dashed line), After this transition, it becomes higher in energy than the smaller total spin states and is no longer the ground state. Each $s_{z}$ GS other than the $s_{z}=4$ GS has a different total spin below and above the transition.
@@
Figure 3: Ground state correlation function for the $3\times 3$ system below and above the NF transition when the lower-left array-site has a spin-up electron. The correlation function for each non-negative spin-$z$ ($s_{z}=3,2,1,0$) ground state is shown. The negative $s_{z}$ GSs have the same plots but for correlation with a spin-down instead of a spin-up electron. The correlation function ranges from -1 to 1, indicated as the color ranges from dark red to dark blue. The arrow points in the upper right direction when the site is positively correlated, and to the bottom left when it is negatively correlated. In the NF regime, below the transition point, the ground state is ferromagnetic; above the transition, the ground state has a smaller total spin, the correlation plots show quasi-antiferromagnetic characteristics.
@@
In the $2\times 2$ array geometry the four sites are equivalent. The $3\times 3$ array geometry is more complicated than the $2\times 2$ geometry because there are three different, inequivalent kinds of sites: the four corner sites are like the $2\times 2$ sites with two connections only to other external sites; the four edge sites that have one additional hopping to the center site, and one center site that is the only internal site, fully connected with four nearest neighbor hoppings. Despite the more complicated structure, the energies for the $s_{z}$ GSs for the $3\times 3$ array follows a similar pattern to the GS energies of the $2\times 2$ array, as can seen by comparing Figs. 1b and 2. The transition is at $|t/U|\approx 0.0144$, which is lower than the transition for the $2\times 2$ array at $0.053$, showing that larger interaction is needed for the transition to occur in the $3\times 3$ array.
@@
where $\sigma_{o,\uparrow/\downarrow}$ indicates the spin up/down occupation of site $o$, $\alpha[\sigma_{f}]$ is the probability amplitude for any configuration of site occupations that includes the $\sigma_{f}$ occupation, and $\alpha[\sigma_{f},\sigma_{o,\uparrow/\downarrow}]$ is the probability amplitude for any configuration that includes the site occupations at $f$ and $o$. The amplitudes’ squared are summed over all configurations with the appropriate site occupations. The correlation function of the $s_{z}$ GS for the $3\times 3$ array when one spin-up electron is fixed on the corner is plotted in Fig. 3. The color indicates the strength of the correlation, if the other site is dark blue, the density on that site is has more spin-up than spin-down and is positively correlated with the fixed up-spin at the corner site. A dark red site indicates the site has larger spin-down density. The direction of the arrow at each site also shows whether the site is positively or negatively correlated with the spin at the fixed site. Fig.3 compares the correlation function below and above the transition point for $s_{z}=3,2,1,0$. Correlations that are ferromagnetic below the transition becomes locally antiferromagnetic above the transition for the small $s_{z}$ GSs.
Figure 4: Correlation function for energies for $s_{z}=3$ as a function of $|t/U|$ near the transition point, offset by the maximum-spin-state energy. A spin-up electron is fixed at the lower-left array-site. Correlation plots are displayed at markers. Green triangles indicate a ferromagnetic state, while red circles indicate a state with same $s_{z}=3$ but smaller $S$. The black dotted line is the transition point. Note that for $s_{z}=3$ subspace, the transition point is larger at $|t/U|\approx 0.0157$ instead of $0.0144$. Markers are $t/U$ values chosen arbitrarily to show how correlation plots change ($|t/U|=0.0125,0.0153,0.0162,0.018$).
From the correlation plots shown in Figs. 3 and 4, the states below and above the transition look qualitatively different because the GS for each spin $z$ switches to a state with the same spin $z$ but smaller total spin $S$. This is confirmed by following the energies for the first few excited states, which have a different total spin from the ground state. Fig. 4 is shows a blow-up of the region near the transition. Below the transition, the ground state energy is the maximum-total-spin state. An excited state below the transition with the same $s_{z}$ but smaller total $S$ becomes the ground state above the transition. As $|t/U|$ moves away from the transition point, the correlation functions for different total-spin states do not change qualitatively, but vary slowly and continuously across and away from the transition.

### v2

@@ Abstract
Nagaoka ferromagnetism (NF) is a long-predicted example of itinerant ferromagnetism (IF) in the Hubbard model that has been studied theoretically for many years. The condition for NF, an infinite on-site Coulomb repulsion and a single hole in a half-filled band, does not arise naturally in materials. NF was only realized recently for the first time in experiments on a $2\times 2$ array of gated quantum dots. Dopant arrays and gated quantum dots in Si allow for engineering controllable systems with complex geometries. This makes dopant and quantum dot arrays good candidates to study NF in different array geometries through analog quantum simulation. Here we present theoretical simulations done for $3\times 3$ arrays and larger $N\times N$ arrays, and predict the emergence of different forms of ferromagnetism in different geometries. We find NF in perfect $3\times 3$ arrays, as well as in $N\times N$ arrays for one hole doping of a half-filled band. The ratio of the hopping $t$ to Hubbard on-site repulsion $U$ that defines the onset of NF scales as $1/N^{3.5}$ as $N$ increases, approaching the bulk limit of infinite $U$ for large $N$. Additional simulations are done for geometries made by removing sites from $N\times N$ arrays. Different forms of ferromagnetism are found for different geometries. Loops show ferromagnetism, but only for three electrons. For loops, the critical $t/U$ for the onset of ferromagnetism scales as $N$ as the loop length increases. We show that the different dependences on size for loops and $N\times N$ arrays can be understood by scaling arguments that highlight the different energy contributions to each different form of ferromagnetism. Our results show how analog quantum simulation with small arrays can elucidate the role of effects including wavefunction connectivity; system geometry, size and symmetry; bulk and edge sites; and kinetic energy in determining quantum magnetism of small systems.
@@ I Introduction
NF is one of the exact results for itinerant ferromagnetism (IF) and has been studied theoretically for years in various conditions [2, 3]. However, the specific conditions proved by Nagaoka do not arise in natural materials and was observed experimentally for the first time in an engineered $2\times 2$ quantum dot plaquette [4]. It remains an open question when NF can be extended to larger, finite-systems that have internal as well as edge sites, how NF would behave in realistic conditions where $U$ is large but not infinite, and how NF is affected by long-range interactions or disorder.
Following Nagaoka’s original work, many have tried to generalize Nagaoka’s result. Tasaki provided a proof of a generalized version of Nagaoka’s theorem [2]. Mielke and others found a new class of ferromagnetic states in the Hubbard model which they called “flat-band ferromagnetism” [5]. The possibility of extending NF to finite $U$ with finite densities of holes has also been studied for different infinite lattices [6, 7]. There have also been many theoretical works done on the instability of NF, that show when NF does not take place [8, 9, 10]. Recently, additional theoretical work has elucidated the polaronic correlation around holes in NF [11, 12].
@@
Quantum dots, also known as artificial atoms, can be used to perform simulations beyond what classical computers can do [17, 18, 19, 20]. Among various experimental platforms, dopant-based quantum dots have many advantages, such as the precise placement of dopant atoms and tunable site-to-site hopping, making them a strong candidate for simulating strongly correlated systems, like the Hubbard model in the large $U$ limit. Silver recently fabricated $3\times 3$ arrays of single/few-dopant quantum dots and demonstrated analog quantum simulation of a 2D Fermi-Hubbard model with tunable parameters [21]. This was the first two-dimensional quantum-dot simulation with internal sites. This provides an experimental pathway to probe NF in $3\times 3$ arrays and in many other finite geometries. Gated semiconductor quantum dots show similar promise. The original effort to study NF in a $2\times 2$ plaquette [4] has been extended to study exciton and spin dynamics in $2\times 4$ ladders. [22, 23] Linear chains formed from multidopant quantum dot clusters has been used to study topological states in the Su-Schrieffer-Heeger model. [24]
@@
In this paper, we simulate $3\times 3$ dopant arrays in the conditions required for NF, i.e. one less electron than a half-filled band and in the large $U$ limit. We predict that NF will exist in a $3\times 3$ array, as in a $2\times 2$ array, as well as in $N\times N$ arrays, for a finite region of $t/U$, where $t$ is the nearest-neighbor hopping. For $t/U$ below the transition point, the ground state of the system has maximum total spin $S$, i.e. the ground state is ferromagnetic. Since the direction of the ferromagnetism is arbitrary, the system energy only depends on total spin $S$ but not the spin in the $z$-direction ($s_{z}$). The ferromagnetic state with maximum $s_{z}=S$ with all of the spins aligned in the same $z$ direction is degenerate with all other ferromagnetic states with smaller $s_{z}$ as long as they have the same total $S$. The ground state is $(2S+1)$-fold degenerate when there is Nagaoka ferromagnetism. Above the transition point, the ground state is no longer the maximum-total-spin state and is no longer ferromagnetic. We also simulate variations of the perfect $3\times 3$ array to determine when NF is robust to experimental impurities or disorder.
@@
The quantum nature of magnetism plays an essential role in determining how and when ferromagnetism occurs. The small-scale simulators that we consider in this paper provide an excellent testbed for investigating quantum effects. We find that the connectivity of the many-body wavefunction plays a key role in determining when NF can occur. Wavefunction connectivity also determines when forms of itinerant ferromagnetism different from NF at fillings other than one hole in a half-filled band can occur. Band filling is another key condition that determines the type of ferromagnetism. We will show that very different fillings, related to the number of holes in a half-filled band or the number of electrons in an empty band, will define the possibility of ferromagnetism in similar, related structures. The competition between kinetic energy and coulomb repulsion effects determines where in $t/U$ space the transition to ferromagnetism occurs. The $t/U$ ratio at the transition can increase or decrease with system size, depending on the array geometry (and the related wavefunction connectivity), and the band filling. We develop simple scaling arguments to explain both types of size-dependence and the critical role that the kinetic energy plays, through its dependence on band filling, in determining the form of the scaling. The role of these quantum effects will be highlighted as we discuss NF in different small arrays to show how quantum simulation with small arrays could reveal these critical effects. It should be emphasized that dopant arrays, once they can be made perfectly [25] would be a versatile solid-state platform to explore the diversity of arrays and the range of effects that we discuss. Gated quantum dots will also be a versatile solid-state platform as techniques to control dots in complex geometries are developed.
@@
The paper is structured as follows: In Section II, we discuss the 2D Hubbard model, the signatures of finding NF in finite systems, and how the systems are modelled in our simulations. In Section III, we present our simulation results for the $3\times 3$ and $N\times N$ square arrays, which exhibit NF up to a finite $t/U$ transition point that decreases with increasing $N$. In Section IV, we describe the robustness of NF in a $3\times 3$ array with disordered sites, that is sites that have been removed. Most importantly, we show that a loop array, i.e. an $N\times N$ with the internal sites removed, possesses a ferromagnetic ground state for three-electron filling with a $t/U$ transition that scales linearly with loop length $N$, in stark contrast to NF in perfect, square arrays. In this section, we also present results to show how the $t/U$ ratio at the transition to ferromagnetism scales with system size $N$. Scaling arguments explain this dependence on $N$ and elucidate the competition between kinetic energy and repulsive interaction affects to determines when the transition occurs. Finally in Section V, we summarize and discuss essential results about quantum magnetism that can be revealed by simulations done on small arrays of dopants or quantum dots.
@@ II Hubbard Model
where $t$ is the hopping between nearest neighbor sites, $c^{{\dagger}}$ and $c$ are the creation and annihilation operators, $n$ is the particle number operator, and $U$ is the on-site interaction between spin-up and spin-down electrons. NF is determined by a competition between hopping energy and repulsion, hence the parameter that matters is $t/U$. Some use the hopping term $t$ with a minus sign, but here we use the convention without a minus sign and allow the value of $t$ to be positive or negative. The sign of $t$ is important because in some systems the Nagaoka theorem can be proved for one sign but not the other. The sign of t can often be changed with a basis change. In particular, the transformation between the signs of $t$ can be done for a bipartite system. In these cases, NF, if it occurs, will occur for both signs of $t$. The $N\times N$ arrays we study are bipartite lattices.
To find out if a system has NF, it is necessary to determine whether the one-hole ground state has maximum total spin. For a finite system like the $2\times 2$ array with 3 electrons, the Hamiltonian is small and the energy spectrum can be calculated directly, as shown in [4]. The energies for the two different total spins, $S=3/2$ and $S=1/2$ are plotted against the parameter $t/U$ in Fig. 1a. Both energies are shifted with respect to the energy of the maximum-total-spin state $E_{S=3/2}=-2t$. Below $|t/U|\approx 0.053$, the ground state of the system is the maximum-total-spin $S_{max}=3/2$ state. In the ferromagnetic ground state, the states with $S=S_{max}$ have a degeneracy of $12S_{max}+$ 1. In the $2\times 2$ array with 3 electrons, there are four degenerate $s_{z}$ states with $s_{z}=-3/2,-1/2,1/2,3/2$. Because the Hubbard model does not include spin-mixing, ED can be done for each $s_{z}$ subspace, thereby reducing the size of each diagonalization. In our simulation, instead of plotting the ground state energy for each total spin $S$, we plot the ground state energy for the different spin-$z$ subspaces as shown in Fig. 1b, and shift them by the energy of the state with maximum $s_{z}$ $E_{s_{z}=3/2}=-2t$. Below the transition when the system is in the NF regime, the ground state is $4$-fold degenerate. The signature of NF is when the ground states of the spin-$z$ subspaces become degenerate with each other. This guarantees that the ground state of the system has maximum total spin.
@@
Figure 1: Nagaoka ferromagnetism in a $2\times 2$ array. (a) Energy spectrum ($S=3/2$ and $S=1/2$ as a function of $|t/U|$), which are offset by the energy of $S=3/2$. Black dotted line indicates the transition for NF. (b) Ground state for each of the different spin z subspaces as a function of $|t/U|$, which are offset by the energy for $s_{z}=3/2$. Above and below the transition point, the ground state for $s_{z}=\pm 1/2$ switches from total $S=3/2$ to $S=1/2$. NF is shown by the shaded region.
Fig. 1 shows how energy behaves as $|t/U|$ varies. The lattice is bipartite so the sign of $t$ does not matter here, and $t$ and $-t$ give the same energies. This is not always the case as seen later when the system is no longer bipartite, and the ±$\pm t$ cases can behave very differently. In Fig. 1b, the $s_{z}=\pm 1/2$ state is always a ground state, but the ground state after the transition has a different total $S$ and degeneracy compared to before the transition. This change in total spin can be verified by direct calculation of the total spin for each spin $z$.
The existence of NF in $2\times 2$ array has been verified experimentally in [4] for a $2\times 2$ gated quantum dot plaquette. An ab initio calculation of NF in quantum dots was done in [26]. Its robustness was discussed in [26] and theoretically in [15].
@@ III.1 $3\times 3$ case
The $2\times 2$ array has no internal sites. However, the $3\times 3$ dopant array studied by [21] has one internal site. In this section, we look at $3\times 3$ and larger arrays to see how NF behaves as the size of the system increases and the system approaches the 2D bulk limit. NF has not yet been presented for the largest finite systems discussed here. The largest arrays consider here are $8\times 8$ arrays, which should be accessible with dopant arrays. These arrays have more internal sites than edge sites and should represent to transition to the 2D bulk limit. Our results will define the $t/U$ ratios needed for NF. For the $3\times 3$ array, the ground state energy for each spin-$z$ subspace is calculated to determine the presence of NF, and if it exists, to identify the point at which the transition occurs. The half-filled band with one hole has $8$ electrons, and the maximum total spin is $4$. If all $9s_{z}$ subspaces have the same ground-state energy, the system has NF.
@@
Figure 2: NF in a $3\times 3$ array. Black dotted line indicates the transition for NF. Ground state for each of the different spin $z$ subspaces plotted as a function of $|t/U|$, which are offset by the energy for $s_{z}=4$. When the plotted energy difference for an $s_{z}$ ground state becomes negative, the state switches from maximum total spin $S=4$ to a lower total spin $S$.
@@
We plot the ground state (GS) energy for each subspace in Fig. 2 (again shifted by the GS energy for maximum $s_{z}$). For a small $t/U$ or a large $U$, the $s_{z}$ states are all degenerate, i.e. the system has a maximum-total-spin ground state. This is in contrast to the usual non-ferromagnetic spin order near half-filling, where states have increasing energy as $s_{z}$ increases, as shown in Fig. 2. The maximum spin GS holds up to $|t/U|\approx 0.0144$ (black dashed line). After this transition, it becomes higher in energy than the smaller total spin states and is no longer the ground state. Each $s_{z}$ GS other than the $s_{z}=4$ GS has a different total spin below and above the transition.
@@
Figure 3: Ground state correlation function for the $3\times 3$ system below and above the NF transition when the lower-left array-site has a fixed spin-up electron. The correlation function for each non-negative spin-$z$ ($s_{z}=3,2,1,0$) ground state is shown. The negative $s_{z}$ GSs have the same plots but for correlation with a spin-down instead of a spin-up electron. The correlation function ranges from -1 to 1, indicated as the color ranges from dark red to dark blue. The arrows point in the upper right direction when the site is positively correlated, and to the bottom left when it is negatively correlated. In the NF regime, below the transition point, the ground state is ferromagnetic; above the transition, the ground state has a smaller total spin, the correlation plots show quasi-antiferromagnetic characteristics.
@@
In the $2\times 2$ array geometry the four sites are equivalent. The $3\times 3$ array geometry is more complicated than the $2\times 2$ geometry because there are three different, inequivalent kinds of sites: the four corner sites are like the $2\times 2$ sites with two connections only to other external sites; the four edge sites that have one additional hopping to the center site, and one center site that is the only internal site, with four nearest neighbor hoppings. Despite the more complicated structure, the energies for the $s_{z}$ GSs for the $3\times 3$ array follow a similar pattern to the GS energies of the $2\times 2$ array, as can seen by comparing Figs. 1b and 2. The transition is at $|t/U|\approx 0.0144$, which is lower than the transition for the $2\times 2$ array at $0.053$, showing that larger interaction is needed for the transition to occur in the $3\times 3$ array.
@@
where $\sigma_{o,\uparrow/\downarrow}$ indicates the spin up/down occupation of site $o$, $\alpha[\sigma_{f}]$ is the probability amplitude for any configuration of site occupations that includes the $\sigma_{f}$ occupation, and $\alpha[\sigma_{f},\sigma_{o,\uparrow/\downarrow}]$ is the probability amplitude for any configuration that includes the site occupations at $f$ and $o$. The amplitudes’ squared are summed over all configurations with the appropriate site occupations. The correlation function of the $s_{z}$ GS for the $3\times 3$ array when one spin-up electron is fixed on the corner is plotted in Fig. 3. The color indicates the sign and strength of the correlation, dark blue means the density on that site has more spin-up than spin-down and is positively correlated with the fixed up-spin at the corner site. A dark red site indicates the site has larger spin-down density. The direction of the arrow at each site also shows whether the site is positively or negatively correlated with the spin at the fixed site. Fig.3 compares the correlation function below and above the transition point for $s_{z}=3,2,1,0$. Correlations that are ferromagnetic below the transition becomes locally more antiferromagnetic above the transition for all $s_{z}$ spaces.
Figure 4: Correlation function and energies for $s_{z}=3$ as a function of $|t/U|$ near the transition point. The energies are offset by the $s_{z}=4$ state energy. A spin-up electron is fixed at the lower-left array-site. Correlation plots are displayed at marked energies. Green triangles indicate a ferromagnetic state ($s_{z}=3$, $S=4$), while red circles indicate a state with a smaller total spin ($s_{z}=3$, $S=3$). The black dotted line is the transition point. Note that for $s_{z}=3$ subspace, the transition point is larger at $|t/U|\approx 0.0157$ instead of $0.0144$. Markers are $t/U$ values chosen arbitrarily to show how correlation plots change ($|t/U|=0.0125,0.0153,0.0162,0.018$).
From the correlation plots shown in Figs. 3, the states below and above the transition look qualitatively different because the GS for each spin $z$ switches to a state with the same spin $z$ but smaller total spin $S$. This is confirmed by following the first few excited states through transition. Fig. 4 shows a blow-up of the region near the transition. Below the transition, the ground state is the maximum-total-spin state. An excited state below the transition with the same $s_{z}$ but smaller total $S$ becomes the ground state above the transition. As $|t/U|$ moves away from the transition point, the correlation functions for different total-spin states do not change qualitatively, but vary slowly and continuously across and away from the transition.